3. Wat is equilibrant vector2 Draw
Answer #1

Solution:   

The equilibrant vector is a vector that are equal and opposite to the resultant vector . It is used to equilibrium condition for the system .

Suppose two force F1 = 50 N and  F2 = 50 N apply on a system with 450 and 1350 angle respectively (shown in figure). 70172N If= F = 50N 270 F = equilibrant = 7072N vectorThe resultant force is

F = 50 Cos450 + 50 Cos450

F = 100 Cos450

F = 100/1.414

F = 70.72 N

and the resultant angle θ = (θ12) /2 = (45+135)/2

θ = 900

Hence for the equilibrium condition we apply the force F' 70.72 N (equal to the resultant force) with angle 2700 (opposite to the resultant angle) on the system . This applying force vector known as Equilibrant vector .